OpenStudy (anonymous):

Need help with this one find a plylnominal p(x), with zeros -1,3 such that P(4) = 10

OpenStudy (anonymous):

\[reqd. ~polynomial~ is ~P(x)=\left( x+1 \right)\left( x-3 \right)\left( x-4 \right)\] simplify it.

OpenStudy (anonymous):

So the signs switch and p(4) is actually part of the factor and not the degree of the polynomial ?

OpenStudy (anonymous):

if f(x1)=0,then x-x1 is a factor of f(x)

OpenStudy (anonymous):

it is apart of degree of polynomial as well.

OpenStudy (anonymous):

i am wrong ,i thought it is P(4)=0

OpenStudy (anonymous):

there should not be factor of x-4 .it should be P(x)=a(x+1)(x-3)

OpenStudy (anonymous):

find a by using p(4)=10

OpenStudy (johnweldon1993):

Alright for this type of problem ...it is important to remember that *anything times a polynomial will make the polynomial still have the same intercepts Hence...when you are given the zeros of a polynomial...you can write it out as surjithayer has...as f(x) = (x + 1)(x - 3) But you can also write it as f(x) = C(x + 1)(x - 3) Because that 'c' wont change anything in terms of intercepts So you would multiply those out to get f(x) = C(x^2 - 2x - 3) That 'C' will help us solving for the P(4) = 10 So now we plug in 4 for the 'x' and make it equal 10 C(4^2 - 2(4) - 3) = 10 C(16 - 8 - 3) = 10 C(5) = 10 C = 2 Now...if C = 2...we can multiply that through the polynomial we had before f(x) = C(x^2 - 2x - 3) Will now become f(x) = 2(x^2 - 2x - 3) f(x) = 2x^2 - 4x - 6)

OpenStudy (anonymous):

P(4)=a(4+1)(4-3)=10 5a=10 a=10/5=2 P(x)=2(x+1)(x-3)

OpenStudy (anonymous):

Thanks for the help you guys rock!

OpenStudy (anonymous):

you are welcome.Sorry for inconvenience.